# Regret Minimization in Stackelberg Games with Side Information

### v2

@@ 3 On the Impossibility of Fully Adversarial No-Regret Learning
4 Limiting the Power of the Adversary
Definition 4.1 (Contextual Follower Best-Response Region).
Definition 4.2 (Contextual Best-Response Region).
One useful property of all contextual best-response regions is that they are convex and bounded polytopes. To see this, observe that every contextual follower best-response region (and therefore every contextual best-response region) is (1) a subset of $\Delta^{d}$ and (2) the intersection of finitely-many half-spaces.
While every $\mathcal{X}_{\mathbf{z}}(\sigma)$ is convex and bounded, it is not necessarily closed. If every contextual best-response region were closed, it would be without loss of generality for the leader to restrict themselves to the set of policies which map every context to an extreme point of some contextual best-response region. In what follows, we show that the leader does not lose too much (measured by regret) by restricting themselves to policies which map to some approximate extreme point of a contextual best-response region.
Definition 4.3 ($\delta$-approximate extreme points).
Note that Definition 4.3 is constructive. We set $\delta=\mathcal{O}(\frac{1}{T})$ so that the additional regret from only considering policies which map to points in $\cup_{\mathbf{z}\in\mathcal{Z}}\mathcal{E}_{\mathbf{z}}(\delta)$ is negligible. As a result, we use the shorthand $\mathcal{E}_{\mathbf{z}}:=\mathcal{E}_{\mathbf{z}}(\delta)$ throughout the sequel. The following lemma is a generalization of Lemma 4.3 in Balcan et al. [6] to the contextual Stackelberg game setting, and its proof uses similar techniques from convex analysis.
Lemma 4.4.
For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathcal{E})}:\mathcal{Z}\rightarrow\cup_{\mathbf{z}\in\mathcal{Z}}\mathcal{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathcal{E}_{\mathbf{z}}$ and obtains the following performance guarantee:
$\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t})))\leq 1$
Moreover, the same result holds in expectation over any distribution over follower types $\mathcal{F}$.
Since we do not restrict the context space to be finite, the leader cannot pre-compute $\mathcal{E}_{\mathbf{z}}$ for every $\mathbf{z}\in\mathcal{Z}$ before the game begins. Instead, they can compute $\mathcal{E}_{\mathbf{z}_{t}}$ in round $t$ before they commit to their mixed strategy. While $\mathcal{E}_{\mathbf{z}_{t}}$ is computatable, it may be exponentially large in $A_{f}$ and $K$. However this is to be expected as Li et al. [24] show that in its general form, the non-contextual version of the online Stackelberg game problem is NP-Hard.
4.1 Stochastic Follower Types and Adversarial Contexts
Definition 4.5 (Expected Contextual Stackelberg Regret).
Under this setting, the expected utility for any fixed policy $\pi$ may be written as
@@ 4.1 Stochastic Follower Types and Adversarial Contexts
and acting greedily with respect to our estimate. Here $\widehat{p}_{t}(b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t}))=a_{f})$ is the (estimated) probability that the follower’s best-response is $a_{f}$, given context $\mathbf{z}_{t}$ and leader mixed strategy $\pi(\mathbf{z}_{t})$. As we will see, different instantiations of $\widehat{p}_{t}$ may lead to different regret rates for Algorithm 1. However, before instantiating Algorithm 1 with a specific estimation procedure, we provide a general result which bounds the regret of Algorithm 1 in terms of the total variation distance between the sequence $\{\widehat{p}_{t}\}_{t\in[T]}$ and the true distribution $p$.
Input: $\widehat{\mathbf{p}}_{t}$
for $t=1,\ldots,T$ do
      Observe context $\mathbf{z}_{t}$, commit to mixed strategy $\mathbf{x}_{t}=\pi_{t}(\mathbf{z}_{t})=\arg\max_{\mathbf{x}\in\mathcal{E}_{\mathbf{z}_{t}}}\widehat{\mathbb{E}}_{t}[u(\mathbf{z}_{t},\mathbf{x},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x})]$, where $\widehat{\mathbb{E}}_{t}[u(\mathbf{z}_{t},\mathbf{x},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x})]$ is computed as in Equation 5.
      Update $\widehat{\mathbf{p}}_{t}\rightarrow\widehat{\mathbf{p}}_{t+1}$
      end for
ALGORITHM 1 Learning with stochastic follower types: full feedback
Theorem 4.6.
Theorem 4.6 shows that the regret of Algorithm 1 is proportional to how well it estimates $\mathbf{p}(\mathbf{z},\mathbf{x})$ over time (as measured by total variation distance), on (1) the sequence of contexts chosen by the adversary and (2) the sequence of mixed strategies played by Algorithm 1 and the (near-)optimal policy $\pi^{(\mathcal{E})}$. While we instantiate Algorithm 1 in the setting where there are finitely-many follower types and follower actions, Theorem 4.6 opens the door to provide meaningful regret guarantees in settings in which there are infinitely-many follower types and/or follower actions. [^8] All that is required to run Algorithm 1 in a particular problem instance is an oracle for evaluating $\widehat{\mathbb{E}}_{t}[u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))]$ and updating $\widehat{\mathbf{p}}_{t}$.
@@
We now instantiate the estimation procedure in Algorithm 1 in two different ways to get end-to-end regret guarantees. In Corollary 4.7, we show that the leader can get regret $O(K\sqrt{T\log T})$ by estimating the distribution of follower types directly, and in Corollary 4.8 we show that the leader’s regret is bounded as $O(A_{f}\sqrt{T\log T})$ if they instead estimate the probability that the follower best-responds with action $a_{f}\in\mathcal{A}_{f}$, given context a particular context and leader mixed strategy.
Corollary 4.7.
If $\;\widehat{\mathbf{p}}_{t}=\{\widehat{p}_{t}(f_{t}=\alpha^{(i)})\}_{i\in[K]}$, $\widehat{p}_{t+1}(f=\alpha^{(i)})=\frac{1}{t}\sum_{\tau=1}^{t}\mathbbm{1}\{f_{\tau}=\alpha^{(i)}\}$, and $\widehat{p}_{1}(f=\alpha^{(i)})=\frac{1}{K}$ for all $i\in[K]$, then the expected regret of Algorithm 1 satisfies
$\mathbb{E}[R(T)]=O\left(K\sqrt{T\log(T)}\right).$

### v3

@@ 3 On the impossibility of fully adversarial no-regret learning
4 Limiting the power of the adversary
Definition 4.1 (Contextual Follower Best-Response Region).
Definition 4.2 (Contextual Best-Response Region).
One useful property of all contextual best-response regions is that they are convex and bounded polytopes. To see this, observe that every contextual follower best-response region (and therefore every contextual best-response region) is (1) a subset of $\Delta^{d}$ and (2) the intersection of finitely-many half-spaces. While every $\mathcal{X}_{\mathbf{z}}(\sigma)$ is convex and bounded, it is not necessarily closed. If every contextual best-response region were closed, it would be without loss of generality for the leader to restrict themselves to the set of policies which map every context to an extreme point of some contextual best-response region. In what follows, we show that the leader does not “lose too much” (as measured by regret) by restricting themselves to policies which map to some approximate extreme point of a contextual best-response region.
Definition 4.3 ($\delta$-approximate extreme points).
Note that Definition 4.3 is constructive. We set $\delta=\mathcal{O}(\frac{1}{T})$ so that the additional regret from only considering policies which map to points in $\cup_{\mathbf{z}\in\mathcal{Z}}\mathcal{E}_{\mathbf{z}}(\delta)$ is negligible. As a result, we use the shorthand $\mathcal{E}_{\mathbf{z}}:=\mathcal{E}_{\mathbf{z}}(\delta)$ throughout the sequel. The following lemma is a generalization of Lemma 4.3 in Balcan et al. [5] to our setting, and its proof uses similar techniques from convex analysis.
Lemma 4.4.
For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathcal{E})}:\mathcal{Z}\rightarrow\cup_{\mathbf{z}\in\mathcal{Z}}\mathcal{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathcal{E}_{\mathbf{z}}$ and guarantees that $\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t})))\leq 1$. Moreover, the same result holds in expectation over any distribution over follower types $\mathcal{F}$.
Since we do not restrict the context space to be finite, the leader cannot pre-compute $\mathcal{E}_{\mathbf{z}}$ for every $\mathbf{z}\in\mathcal{Z}$ before the game begins. Instead, they can compute $\mathcal{E}_{\mathbf{z}_{t}}$ in round $t$ before they commit to their mixed strategy. While $\mathcal{E}_{\mathbf{z}_{t}}$ is computatable, it may be exponentially large in $A_{f}$ and $K$. However this is to be expected as Li et al. [23] show that in its general form, solving the non-contextual version of the online Stackelberg game problem is NP-Hard.
4.1 Stochastic follower types and adversarial contexts
Definition 4.5 (Expected Contextual Stackelberg Regret).
Under this setting, the utility for policy $\pi$ may be written as
[…]
and acting greedily with respect to our estimate. Here $\widehat{p}_{t}(b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t}))=a_{f})$ is the (estimated) probability that the follower’s best-response is $a_{f}$, given context $\mathbf{z}_{t}$ and leader mixed strategy $\pi(\mathbf{z}_{t})$. As we will see, different instantiations of $\widehat{p}_{t}$ will lead to different regret rates for Algorithm 1. However, before instantiating Algorithm 1 with a specific estimation procedure, we provide a general result which bounds the regret of Algorithm 1 in terms of the total variation distance between the sequence $\{\widehat{p}_{t}\}_{t\in[T]}$ and the true distribution $p$.
Input: $\widehat{\mathbf{p}}_{t}$
for $t=1,\ldots,T$ do
      Observe $\mathbf{z}_{t}$, commit to $\mathbf{x}_{t}=\pi_{t}(\mathbf{z}_{t})=\arg\max_{\mathbf{x}\in\mathcal{E}_{\mathbf{z}_{t}}}\widehat{\mathbb{E}}_{t}[u(\mathbf{z}_{t},\mathbf{x},b_{f_{t}}(\mathbf{z}_{t},\mathbf{x})]$ (Equation 1).
      Update $\widehat{\mathbf{p}}_{t}\rightarrow\widehat{\mathbf{p}}_{t+1}$
      end for
Algorithm 1 Learning with stochastic follower types: full feedback
Theorem 4.6.
Theorem 4.6 shows that the regret of Algorithm 1 is proportional to how well it estimates $\mathbf{p}(\mathbf{z},\mathbf{x})$ over time (as measured by total variation distance), on (1) the sequence of contexts chosen by the adversary and (2) the sequence of mixed strategies played by Algorithm 1 and the (near-)optimal policy $\pi^{(\mathcal{E})}$. While we instantiate Algorithm 1 in the setting where there are finitely-many follower types and follower actions, Theorem 4.6 opens the door to provide meaningful regret guarantees in settings in which there are infinitely-many follower types and/or follower actions. [^6] All that is required to run Algorithm 1 in a particular problem instance is an oracle for evaluating $\widehat{\mathbb{E}}_{t}[u(\mathbf{z},\mathbf{x},b_{f_{t}}(\mathbf{z},\mathbf{x}))]$ and updating $\widehat{\mathbf{p}}_{t}$. We now instantiate the estimation procedure in Algorithm 1 in two different ways to get end-to-end regret guarantees. First, the leader can get regret $O(K\sqrt{T\log T})$ by estimating the distribution of follower types directly.
Corollary 4.7.
If $\;\widehat{\mathbf{p}}_{t}=\{\widehat{p}_{t}(f_{t}=\alpha^{(i)})\}_{i\in[K]}$, $\widehat{p}_{t+1}(f=\alpha^{(i)})=\frac{1}{t}\sum_{\tau=1}^{t}\mathbbm{1}\{f_{\tau}=\alpha^{(i)}\}$, and $\widehat{p}_{1}(f=\alpha^{(i)})=\frac{1}{K}$ for $i\in[K]$, then the regret of Algorithm 1 satisfies $\mathbb{E}[R(T)]=O(K\sqrt{T\log(T)})$.
The leader can obtain a complementary regret bound of $O(A_{f}\sqrt{T\log T})$ if they instead estimate the probability that the follower best-responds with action $a_{f}\in\mathcal{A}_{f}$, given a particular context $\mathbf{z}$ and leader mixed strategy $\mathbf{x}$ [^7] In general $K$ has no dependence on $A_{f}$, and vice versa. In what follows, we use $\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a_{f})}\in\{0,1\}^{K}$ to refer to the indicator vector whose $i$-th component is $\mathbbm{1}\{\sigma^{(\mathbf{z},\mathbf{x})}(\alpha^{(i)})=a_{f}\}$, i.e. the indicator that a follower of type $\alpha^{(i)}$ best-responds to context $\mathbf{z}$ and mixed strategy $\mathbf{x}$ by playing action $a_{f}$.
Corollary 4.8.
If $\widehat{\mathbf{p}}_{t}(\mathbf{z},\mathbf{x})=\{\widehat{p}_{t}(\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a_{f})})\}_{a_{f}\in\mathcal{A}_{f}}$, $\widehat{p}_{t+1}(\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a)})=\frac{1}{t}\sum_{\tau=1}^{t}\mathbbm{1}\{b_{f_{\tau}}(\mathbf{z},\mathbf{x})=a\}$, and $\widehat{p}_{1}(\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a)})=\frac{1}{A_{f}}$ for $a_{f}\in\mathcal{A}_{f}$, then the regret of Algorithm 1 satisfies $\mathbb{E}[R(T)]=O(A_{f}\sqrt{T\log(T)})$.
4.2 Stochastic contexts and adversarial follower types
Definition 4.9 (Expected Contextual Stackelberg Regret, II).
Input: Set of weights $\Omega$
[…]
for $t=1,\ldots,T$ do
      Sample $\pi_{t}\sim\mathbf{p}_{t}$, $a_{l,t}\sim\pi_{t}(\mathbf{z}_{t})$, receive utility $u(\mathbf{z}_{t},a_{l,t},b_{f_{t}}(\pi_{t}(\mathbf{z}_{t})))$, observe type $f_{t}$.
[…]
      end for
Algorithm 2 Learning with stochastic contexts: full feedback
Lemma 4.10.
Theorem 4.11.
If $\Omega=\{\boldsymbol{\omega}\;:\;\boldsymbol{\omega}\in\Delta^{K},T\cdot\boldsymbol{\omega}[i]\in\mathbb{N},\forall i\in[K]\}$ and $\eta=\sqrt{\frac{\log\Pi}{T}}$, then Algorithm 2 obtains expected contextual Stackelberg regret (Definition 4.9) $\mathbb{E}[R(T)]=O\left(\sqrt{KT\log T}+K\right)$.
5 Extension to bandit feedback
We have so far assumed that the leader gets to observe the follower’s type after each round. However this assumption may not always hold in real-world Stackelberg game settings. For example, in cyber security domains it may be hard to deduce the organization responsible for a failed cyber attack. In wildlife protection, a very successful poacher may never be seen by the park rangers. Instead, the leader may only be able to observe the action the follower takes at each round. Following previous work on learning in non-contextual Stackelberg games, we refer to this type of feedback as bandit feedback. What can we say about the leader’s ability to learn under bandit feedback when there is side information?
While our impossibility result of Section 3 immediately applies to this more challenging setting, our algorithms from Section 4 do not. This is because we can no longer compute quantities such as $\mathbbm{1}\{f_{t}=\alpha^{(i)}\}$ or $b_{f_{t}}(\mathbf{z}_{t},\mathbf{x})$ for an arbitrary mixed strategy $\mathbf{x}$ from just follower $f_{t}$’s action alone. We still assume that the follower is one of $K$ different types, although $f_{t}$ is now never revealed to the leader.
We allow ourselves two relaxations when designing learning algorithms which operate under bandit feedback. First, while the leader’s utility function may still depend on the context $\mathbf{z}_{t}$, we assume that the follower’s utility is a function of the leader’s mixed strategy $\mathbf{x}_{t}$ alone, i.e. $u_{f}(\mathbf{z},\mathbf{x},a_{f})=u_{f}(\mathbf{x},a_{f})$ for all $\mathbf{z}\in\mathcal{Z}$. This allows us to drop the dependence on $\mathbf{z}_{t}$ from both the follower’s best response and the set of approximate extreme points, i.e. $b_{f}(\mathbf{z},\mathbf{x})$ becomes $b_{f}(\mathbf{x})$ and $\mathcal{E}_{\mathbf{z}}$ becomes $\mathcal{E}$. Furthermore, our definitions of contextual follower best-response region (Definition 4.1) and contextual best-response region (Definition 4.2) collapse to their non-contextual counterparts. Depending on the application domain, the assumption that only the leader’s utility depends on the side information may be reasonable. For instance, while an institution would prefer that a server with less traffic is hacked compared to one with more, a hacker might only care about the information hosted on the server (which may not be related to network traffic patterns). Second, we design algorithms with regret guarantees which only hold against a non-adaptive adversary. [^10] We hypothesize that our results in this section could be extended to hold against an adaptive adversary by using more clever exploration strategies. Despite these relaxations, the problem of learning under bandit feedback still remains challenging because of the exponentially large size of $\mathcal{E}$. While a natural first step is to estimate $p(b_{f_{t}}(\mathbf{x})=a_{f})$ (i.e. the probability that follower at round $t$ best-responds with action $a_{f}$ when the leader plays mixed strategy $\mathbf{x}$) for all $\mathbf{x}\in\mathcal{E}$ and $a_{f}\in\mathcal{A}_{f}$, doing so naively would take exponentially-many rounds, due to the size of $\mathcal{E}$.
Building off of results in the non-contextual setting of Balcan et al. [5], we leverage the fact that the leader’s utility for different mixed strategies is not independent. Instead, they are linearly related through the frequency of follower types which take a particular action, given a particular leader mixed strategy. Therefore, it suffices to estimate this linear function (which can be done using as few as $K$ samples) to get an unbiased estimate of $p(b_{f_{t}}(\mathbf{x})=a_{f})$ for any $\mathbf{x}\in\mathcal{E}$ and $a_{f}\in\mathcal{A}_{f}$. Borrowing from the literature on linear bandits, we use a barycentric spanner [4] to estimate $\{\{p(b_{f_{t}}(\mathbf{x})=a_{f})\}_{a_{f}\in\mathcal{A}_{f}}\}_{\mathbf{x}\in\mathcal{E}}$ in both partial adversarial settings we consider. A barycentric spanner for compact vector space $\mathcal{W}$ is a special basis such that any vector in $\mathcal{W}$ may be expressed as a linear combination of elements in the basis, with each linear coefficient being in the range $[-1,1]$.
In Section C.1, we use the property that estimators constructed using barycentric spanners have low variance to show that an explore-then-exploit algorithm achieves $O(K^{2/3}A_{f}^{2/3}T^{2/3}\log^{1/3}T)$ expected contextual Stackelberg regret in the setting with stochastic follower types and adversarial contexts. Specifically, our algorithm (Algorithm 3) plays a special set of $K$ mixed strategies $N$ times each, then uses barycentric spanners to estimate $\{p_{t}(\mathbbm{1}_{(\sigma^{(\mathbf{z},\mathbf{x})}=a_{f})})\}_{a_{f}\in\mathcal{A}_{f}}$ for all $\mathbf{x}\in\mathcal{X}$ and $\mathbf{z}\in\mathcal{Z}$, after which Algorithm 3 plays greedily like in Section 4.1.
In Section C.2, we use the property that estimators constructed using barycentric spanners are bounded to design a reduction to our algorithm in Section 4.2 which achieves $O(KA_{f}^{1/3}T^{2/3}\log^{1/3}T)$ expected contextual Stackelberg regret whenever the sequence of contexts is chosen stochastically and the sequence of follower types is chosen by an adversary. Finally, while it may be possible to obtain $O(\sqrt{T})$ regret without using barycentric spanners, this would come at the cost of a linear dependence on $|\mathcal{E}|$ (and therefore an exponential dependence on $K$ and $A_{f}$) in regret.
6 Conclusion
We initiate the study of Stackelberg games with side information, which despite the presence of side information in many Stackelberg game settings, has not received attention from the community. We focus on the online setting in which the leader faces a sequence of contexts and follower types. We show that when both sequences are chosen adversarially, no-regret learning is not possible even for highly structured policy classes. When either sequence is chosen stochastically, we obtain algorithms with $\tilde{O}(\sqrt{T})$ regret. We also explore an extension to bandit feedback, in which we obtain $\tilde{O}(T^{2/3})$ regret in both settings. There are several exciting avenues for future research; we highlight two below.
  1. 1.
    Intermediate forms of adversary. The two relaxations of the fully adversarial setting that we consider, while natural, rule out the leader learning about the follower’s type from the context. Although we prove that learning is impossible in the fully adversarial setting, our lower bound does not rule out, e.g. settings where the mapping from contexts to follower types has finite Littlestone dimension. It would be interesting to further explore this direction to pin down when no-regret learning is possible.
  2. 2.
    $\tilde{O}(T^{1/2})$ regret under bandit feedback. Bernasconi et al. [6] obtain $O(T^{1/2})$ regret when learning in non-contextual Stackelberg games under bandit feedback against an adversarially-chosen sequence of follower types via a reduction to adversarial linear bandits. However, applying similar steps to Bernasconi et al. in our setting results in a reduction to a generalization of the (adversarial) contextual bandit problem for which we are not aware of any regret minimizing algorithm. Nevertheless, we view exploring whether $\tilde{O}(T^{1/2})$ contextual Stackelberg regret is possible under bandit feedback as a natural and exciting future direction.
Appendix A Appendix for Section 3: On the impossibility of fully adversarial no-regret learning
See 3.2
@@ Appendix A Appendix for Section 3: On the impossibility of fully adversarial no-regret learning
Appendix B Appendix for Section 4: Limiting the power of the adversary
See 4.4
Proof.
Observe that for any $\mathbf{z}$,
@@ Appendix B Appendix for Section 4: Limiting the power of the adversary
B.1 Section 4.1: Stochastic follower types and adversarial contexts
See 4.6
@@
See 4.7